OpenStudy (swissgirl):

Determine whether the relation O defined on ℤ is reflexive, symmetric, or transitive. The relation O on ℤ is defined as follows: for all m, n ∈ ℤ, m O n ⇔ m – n is odd

OpenStudy (kinggeorge):

Do you know what it means to be reflexive?

OpenStudy (swissgirl):

when there is (x,x) is an element in Z

OpenStudy (swissgirl):

whoops that was not proper english and didnt make sense but u got the jist of things

OpenStudy (kinggeorge):

That's exactly what it means to be reflexive. So now, look at \(xOx\iff x-x\) is odd. Is x-x odd?

OpenStudy (swissgirl):

NOOOOOOOOOOOOO

OpenStudy (swissgirl):

since it will be 0 and 0 isnt odd

OpenStudy (kinggeorge):

Perfect. So that means that it's not reflexive.

OpenStudy (swissgirl):

ummm well i think its syymetric

OpenStudy (kinggeorge):

Onto symmetric. Suppose \(x-y\) is odd. Is \(y-x\) also odd?

OpenStudy (swissgirl):

yesssss

OpenStudy (kinggeorge):

It sure is. So now we only need to look at transitive.

OpenStudy (swissgirl):

hmmmmm that one is a lil more difficult to figure out give me a sec

OpenStudy (swissgirl):

its not transitive

OpenStudy (swissgirl):

x=4 y=3 z=2

OpenStudy (swissgirl):

4-3=1 3-2=1 4-2=2

OpenStudy (kinggeorge):

That's correct. And a counterexample to go along with it too.

OpenStudy (kinggeorge):

In fact, you could prove that it's intransitive by saying that since \(x-y\) is odd, and \(y-z\) is odd, then \((x-y)+(y-z)=(x-z)\) must be even since odd+odd=even.

OpenStudy (kinggeorge):

Rather, that would be antitransitivity, and not intransitivity.

OpenStudy (swissgirl):

huh? lol let me reread it

OpenStudy (swissgirl):

ohhhh i get it

OpenStudy (swissgirl):

YAY U R awesome KING GEORGE

OpenStudy (kinggeorge):

Thank you :)

OpenStudy (kinggeorge):

Good job figuring out most of this by yourself.